OpenStudy (marcelie):

a particle moves along a line so that its velocity at time t is v(t)= t^2-t-6. ( Measured in meters per second ) a. find the displacement of the particle during the time period 1

OpenStudy (marcelie):

@zepdrix

OpenStudy (marcelie):

@freckles

OpenStudy (marcelie):

@satellite73

satellite73 (satellite73):

first one is the integral

satellite73 (satellite73):

\[\int_1^4 (t^2-t-6)dt\]

satellite73 (satellite73):

you good with that?

OpenStudy (marcelie):

yes

satellite73 (satellite73):

second one is not the integral, you have to break it in two parts

OpenStudy (marcelie):

does it need absolute value ? for the one you wrote

satellite73 (satellite73):

no

satellite73 (satellite73):

first one is displacement, means how far you were from where you started

satellite73 (satellite73):

second one is total distance including going forward and backward

OpenStudy (marcelie):

so then the distance does require the absolute value ?

satellite73 (satellite73):

\[t^2-t-6\] is negative until you get to 3, then positive that means the particle is going to the left, then to the right

satellite73 (satellite73):

i guess i should ask first if that is clear or no, and did you see how i knew it

OpenStudy (marcelie):

im sorry im really confuse with these problems i have trouble what to do

satellite73 (satellite73):

ok lets go slow first is done though right? just the integral

OpenStudy (marcelie):

yes so do we integrate that ?

satellite73 (satellite73):

yes, for the first one the displacement is \[\int_1^4 (t^2-t-6)dt\] whatever that is it is an easy enough integral to compute we can do that second if you like, lets to part two first

OpenStudy (marcelie):

okay

satellite73 (satellite73):

\[v(t)=t^2-t-6\] is a parabola that opens up that is clear, yes?

OpenStudy (marcelie):

yes

satellite73 (satellite73):

it also factors nicely as \((t-3)(t+2)\) so it is zero at \(t=-2\) and \(t=3\)

satellite73 (satellite73):

here is the picture http://www.wolframalpha.com/input/?i=t^2-t-6

satellite73 (satellite73):

you can tell even without the picture that is is negative on \((-2,3)\) and positive on \((3,\infty)\)

satellite73 (satellite73):

of course we are only concerned with the interval \((1,4)\) on \((1,3)\) it is negative that means the particle is moving to the LEFT on \((3,4)\) \(v\) is positive, that means it is moving to the RIGHT

satellite73 (satellite73):

so far so good?

OpenStudy (marcelie):

oh okay is there a way to do it algebraically to know that ?

satellite73 (satellite73):

it is pretty obvious right? a parabola that opens up is negative between the zeros and positive outside them

satellite73 (satellite73):

so you have to break the integral in to two parts \[\int_1^3v(t)dt+\int _3^4v(t)dt\] the first one will be negative, make it positive

OpenStudy (marcelie):

yes do its just integrates and itll be f(b)- f(a)

satellite73 (satellite73):

which one we doing now first or second?

satellite73 (satellite73):

part A is displacement you don't split them part B is the distance travelled, that is the one you split

OpenStudy (marcelie):

oh okay a. is -4.5 and now we doing b lol

satellite73 (satellite73):

you want me to check?

OpenStudy (marcelie):

how would be look like ?